• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name L. Zhu

    Log entry time 15:50:18 on April18,2007

    Entry number 199504

    keyword=day shift summary

    8:00 keep taking production data with 92uA beam.
    ---------
    2618 he3 1.2M
    2619 he3 1.2M
    2630 he3 219K
    ------------

    8:45 stop the beam and change to new kinematics at Q2=29 fm-2:
    --------
    beam energy: E=4.057
    target: he4
    HRSL: E_e=3.861; theta_e=15.35
    HRSR: E_h=0.4912; theta_h=74.286
    --------
    In addition, the techs changed the camera for the target. One can now see the burnt mark on the BeO target clearly. And Jack did some test on right Q1.

    4:00 pass to next shift and keep waiting for beam.